OpenStudy (anonymous):

x2/6>7/6 how do i make the variable alone???

OpenStudy (anonymous):

@iambatman batman help please

OpenStudy (anonymous):

\[\large x ^{\frac{ 2 }{ 6 }}>\frac{ 7 }{ 6 }\] first thing is this what you meant?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok well remember 2/6 = 1/3, do you agree?

OpenStudy (anonymous):

yep

geerky42 (geerky42):

\(x^{2/6}\) and not \(\dfrac{x^2}{6}\)? Kind of weird for teacher to give you a problem with non-simplified fraction.

OpenStudy (anonymous):

wait um i think it would be x>14/36 right?

OpenStudy (anonymous):

Ok, so when you deal with such questions, don't be afraid of the sign! Treat it as if it were an = sign, just with some restrictions and what not, so for now we'll just treat as an equal sign as there is no restriction we have to follow (for now). \[x ^{\frac{ 1 }{ 3 }} > \frac{ 7 }{ 6 } \] so how do we cancel 1/3 in the exponent? By cubing both sides. \[x > \left( \frac{ 7 }{ 6 } \right)^3\]

OpenStudy (anonymous):

Can you finish it off?

OpenStudy (anonymous):

@geerky42 the original problem was 1/3x-1/2>2/3

OpenStudy (anonymous):

Mhm

OpenStudy (anonymous):

wait i thought to cancel 1.3 you multiply it by 3?

geerky42 (geerky42):

\[\dfrac{1}{3}x-\dfrac{1}{2}>\dfrac{2}{3}\]?

OpenStudy (anonymous):

1/3*

OpenStudy (anonymous):

Can you draw or use latex for your original problem, kind of confusing.

OpenStudy (anonymous):

yes @geerky42

OpenStudy (anonymous):

Oh ok so how did you get \[x^{2/6}\]

OpenStudy (anonymous):

i made all the numbers with the same numerator.

OpenStudy (anonymous):

Or did I just imagine it wrong?

OpenStudy (anonymous):

i turned the problem to \[2/6 x-3/4>4/6\]

geerky42 (geerky42):

Oh, then you are on right track @chucho78 You are stuck at \(\dfrac{2}{6}x>\dfrac{7}{6}\) To isolate x, start by multiplying both sides by 6. then divide both sides by 2

OpenStudy (anonymous):

Ok that makes sense I thought it was the exponent..

OpenStudy (anonymous):

You need to specify these things and also not say yes when I write it out wrong :P

OpenStudy (anonymous):

thought you where using it as a fraction my bad.

OpenStudy (anonymous):

Np, so yeah multiply both sides by 6/2

OpenStudy (anonymous):

2/6*7/6 thats where im stuck @.@

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

but how would i multiply the right side?

geerky42 (geerky42):

Multipling both sides would get you \(x>\dfrac{7}{6}\cdot\dfrac{6}{2}\)

geerky42 (geerky42):

You can cancel 6 out.

OpenStudy (anonymous):

Thats the confusing part!

OpenStudy (anonymous):